OpenStudy (sighn0more):

Can someone please help me work through this equation? 2(n - 3) + 5 = 3(n - 1)

OpenStudy (bahrom7893):

open up the parenthesis: 2*n + 2*(-3) + 5 = 3*n + 3*(-1)

OpenStudy (bahrom7893):

2n - 2*3 + 5 = 3*n - 3*1 2n - 6 + 5 = 3n - 3

OpenStudy (bahrom7893):

2n - 1 = 3n - 3 move all the variable terms to one side and the constants to the other: 2n - 3n = -3 + 1

OpenStudy (sighn0more):

Okay, I think that I get it now. I can solve the rest. Thank you very much!

OpenStudy (sighn0more):

n = 2

OpenStudy (bahrom7893):

you're right.

OpenStudy (anonymous):

2(n - 3) + 5 = 3(n - 1) 2n-6+5=3n-3 2n-1=3n-3 2n-3n=-3+1 -n=-2 n=2